# Data-Enabled Predictive Control and Guidance for Autonomous Underwater Vehicles

### v1

@@ Abstract
This paper presents a fully data-driven control framework for autonomous underwater vehicles (AUVs) based on Data-Enabled Predictive Control (DeePC). The approach eliminates the need for explicit hydrodynamic modeling by exploiting measured input–output data to predict and optimize future system behavior. Classic DeePC was employed in the heading control, while a cascaded DeePC architecture is proposed for depth regulation, incorporating a loop-frequency separation to handle the different dynamic modes of input and output. For 3-D waypoint path following, the Adaptive Line-of-Sight algorithm is extended to a predictive formulation and integrated with DeePC. All methods are validated in extensive simulation on the REMUS 100 AUV and compared with classical PI/PID control. The results demonstrate superior tracking performance and robustness of DeePC under ocean-current disturbances and nonlinear operating conditions, while significantly reducing modeling effort.
@@
Data-driven Model Predictive Control, Autonomous Underwater Vehicles, Cascaded DeePC, Predictive ALOS, 3-D path following.
thanks: Code available: https://github.com/SebsDevLab/DeePC_AUV.git
@@ 1 Introduction
Autonomous Underwater Vehicles (AUVs) have become central to modern marine exploration, inspection, and defense activities (Gu et al., 2022). Ensuring reliable path-following and trajectory tracking in such missions requires controllers that can cope with nonlinear hydrodynamics and environmental disturbances (Shen and Shi, 2020). These control objectives are typically achieved through a hierarchical control structure, which integrates a high-level guidance system responsible for generating appropriate reference signals, an intermediate-level controller tracking those references, and a low-level control system that directly manages the actuators (Breivik and Fossen, 2009). A classic example of a high-level guidance system for AUVs is Line-of-Sight (LOS) guidance. The LOS method provides a geometric three-point guidance rule for steering the vehicle toward a target point on the desired path, effectively translating the spatial error into a heading command (Breivik and Fossen, 2009; Børhaug and Pettersen, 2006). More sophisticated LOS-based guidance laws have been summarized in (Gu et al., 2022), and the state-of-the-art lies in the Adaptive Line-of-Sight (ALOS) algorithm, which optimizes reference generation by compensating for varying environmental conditions (Fossen, 2023). In practice, these high-level guidance laws are commonly paired with reactive PI/PID controllers that do not account for future behavior or constraints. As a result, their performance deteriorates when faced with strong nonlinearities, actuator saturation, or delays in dynamics. Consequently, predictive control strategies have been introduced to AUVs, such as Model Predictive Control (MPC) (Rout and Subudhi, 2020; Shen and Shi, 2020).
These approaches require a model of the AUV dynamics, which is complex and time-consuming (Taubert et al., 2014). The hydrodynamic coefficients that characterize added mass, drag, lift, and damping are often nonlinear, coupled, and highly dependent on factors such as vehicle geometry, speed, and surrounding flow conditions. Identifying these parameters requires extensive experiments that are time-consuming and context-specific (Taubert et al., 2014), and the resulting models typically remain valid only within a limited operating range, reducing their applicability across different speeds, payloads, or environmental conditions. Such challenges have motivated data-driven control approaches, which exploit measured data to design and execute control laws without requiring an explicit parametric model of the system. Within this context, the Data-Enabled Predictive Control (DeePC) framework introduced in (Coulson et al., 2019) has emerged as one of the most promising approaches. DeePC uses offline trajectories of inputs and outputs to solve a constrained optimization problem for optimal control inputs. This is particularly attractive for AUV applications, especially due to its potential to also handle complex and nonlinear dynamics (Elokda et al., 2021).
@@
This paper is organized as follows. Notation is outlined in Section 2, and the general AUV dynamics, state-of-the-art, and theoretical foundations of DeePC are in Section 3, Section 4, and Section 5, respectively. Subsequently, the control architecture design and PALOS are described in Section 6-7, and numerical results obtained from detailed MSS-based simulations (Fossen and Perez, 2004) are analyzed in Section 8. Additionally, the study includes stress tests addressing system nonlinearities and ocean currents to evaluate the robustness and identify the limitations of the proposed control approaches. The built-in PID controllers and the ALOS algorithm from the MSS toolbox are employed as benchmarks to contextualize the obtained results. Finally, the conclusion and future works are in Section 9.
@@ 3 Autonomous Underwater Vehicle
The AUV considered in this study is the REMUS 100, as implemented in Fossen’s MSS toolbox (Fossen and Perez, 2004). The model includes six degrees of freedom (6-DOF) dynamics with actuation provided by a three-bladed propeller, stern, and rudder fins, allowing control of surge, heave, and sway motions. The REMUS 100 is a representative small vehicle with a weight of approximately 31.9 kg, a length of 1.6 m, and a diameter of 0.19 m. (Fossen, 2021) provides a detailed presentation of the mathematical models and theoretical concepts underlying longitudinal AUVs. In this paper, only the main relations relevant for control design are summarized, while the specific REMUS 100 parameters are obtained from the MSS toolbox (Fossen and Perez, 2004). Appendices B63 provide the explicit dynamic formulations and supplementary material supporting the subsequent derivations.
@@ 3.1 The rigid-body equations of an AUV
The dynamics of the AUV are represented using a standard 6-DOF rigid-body formulation (Fossen, 2021). The vehicle state $x=[\eta^{\top},\nu^{\top}]^{\top}$ is defined by its pose $\eta$ and velocity vector $\nu$. The pose consists of its position $x_{p},y_{p},z_{p}$ in the inertial NED frame and the attitude, consisting of the Euler angles $\phi,\theta,\psi$, which describe the orientation of the body frame relative to the NED frame in (1). The linear velocities $u$, $v$, $w$ (surge, sway, heave) and angular velocities $p$, $q$, $r$ (roll, pitch, yaw) in the velocity vector $\nu$ are expressed in the body frame given in (2). The relative velocity with respect to the surrounding water is given in (3), where $\nu_{c}=[u_{c},v_{c},w_{c},0,0,0]^{\top}$ contains the ocean currents expressed in the body frame.
@@
3.2 AUV dynamic matrices
The detailed derivation of the describing dynamic components of the AUV is attached in Appendix B and C. The kinematic equation, given in (4), is defined by the transformation matrix $J(\eta)$, mapping body-fixed linear and angular velocities to the time derivatives of the vehicle’s position and orientation in the NED frame. The inertia matrix $M$ of the AUV is composed of the rigid-body inertia $M_{RB}$, depending on the vehicle’s mass properties, and the added-mass matrix $M_{A}$, depending on the hydrodynamic contribution caused by the acceleration of the surrounding fluid. In the MSS toolbox, $M_{A}$ is assumed constant, resulting in a constant $M$ as in (6).
${M}=\begin{bmatrix}m_{11}&0&0&0&m_{15}&0\\ 0&m_{22}&0&m_{24}&0&0\\ 0&0&m_{33}&0&0&0\\ 0&m_{42}&0&m_{44}&0&0\\ m_{51}&0&0&0&m_{55}&0\\ 0&0&0&0&0&m_{66}\end{bmatrix}$.
The Coriolis–centripetal matrix $C(\nu_{r})$ in (7) represents gyroscopic and cross-coupling effects arising from linear and angular motion. It is composed of the rigid-body term $C_{RB}(\nu_{r})$, depending on the angular velocities in $\nu_{r}$, and the added-mass term $C_{A}(\nu_{r})$, depending on the relative linear velocities in $\nu_{r}$. Both dependencies contribute to the nonlinear character of the AUV dynamics.
@@ 3.2 AUV dynamic matrices
The damping matrix $D(\nu_{r})$ is modeled as a diagonal matrix, representing damping in each degree of freedom with the damping coefficients $X_{u},Y_{v},Z_{w},K_{p},M_{q},N_{r}$. To account for nonlinear effects at higher speeds, the damping in the surge and sway directions decreases in the REMUS 100 simulation exponentially with the total relative velocity $U_{r}$, introducing velocity-dependent behavior in the damping matrix. The resulting formulation is
$D(\nu_{r})=-\operatorname{diag}\Big\{\,X_{u}e^{-3U_{r}},\;Y_{v}e^{-3U_{r}},\;Z_{w},\;K_{p},\;M_{q},\;N_{r}\,\Big\}.$
@@
${g}({\eta})=\left[\begin{array}[]{c}\mathbf{0}_{3\times 1}\\ z_{\mathrm{g}}mg\cos(\theta)\sin(\phi)-y_{\mathrm{g}}mg\cos(\theta)\cos(\phi)\\ z_{\mathrm{g}}mg\sin(\theta)+x_{\mathrm{g}}mg\cos(\theta)\cos(\phi)\\ x_{\mathrm{g}}mg\cos(\theta)\sin(\phi)-y_{\mathrm{g}}mg\sin(\theta)\end{array}\right]$
@@ 3.3 AUV Forces and Moments
3.4 Ocean current dynamics
Ocean currents are modeled as irrotational, such that only the linear velocity components contribute to the relative motion between the AUV and the surrounding water. The current in the NED frame is decomposed into a horizontal component with magnitude $V_{c}$ and direction $\beta_{c}$, and a vertical component $W_{c}$,
${\nu}_{c}^{n}=\begin{bmatrix}V_{c}\cos(\beta_{c}),\;V_{c}\sin(\beta_{c}),\;W_{c}\end{bmatrix}^{\top}$. (13)
Further in-depth modeling details are provided in Appendix 63.
@@ 4 Classic AUV Control
Since the control objective of this work is path-following, the propeller speed $n_{p}$ is fixed to a constant value corresponding to a nominal cruising velocity. The controller gains of the following PI/PID controllers were tuned by pole placement as described in (Fossen, 2021) and (Beard and McLain, 2012).
@@ 4.2 Depth Control
$\theta_{d}$ $=K_{p_{z}}(z_{p}-z_{p,d})+K_{i_{z}}{\textstyle\int_{0}^{t}(z_{p}-z_{p,d})\,\mathrm{d}\tau},$ (15)
$\delta_{s}$ $=-K_{p_{\theta}}(\theta-\theta_{d})-K_{d_{\theta}}q-K_{i_{\theta}}{\textstyle\int_{0}^{t}(\theta-\theta_{d})\,\mathrm{d}\tau}.$
@@ 4.3 Adaptive Line-of-Sight Guidance Law
The ALOS guidance law (Fossen and Aguiar, 2024) provides a 3-D formulation for path following of AUVs. It generates desired yaw and pitch angles, $\psi_{d}$ and $\theta_{d}$, to align the vehicle with a spatial path while adaptively compensating for unknown constant disturbances such as ocean currents. The spatial path is described by the position vector ${p}_{s}(s)=[x_{s}(s),\,y_{s}(s),\,z_{s}(s)]^{\top}$, parameterized by the path variable $s$ and contains all possible points on the path. This representation can encode smooth curves, splines, or straight lines. Its local orientation is given by the tangent vector ${t}_{t}(s)$, which in this work is simplified to a straight line between two consecutive waypoints as in (16).
@@
$\begin{aligned} \pi_{h}=\arctan\!\left(\frac{y_{t}^{\prime}(s)}{x_{t}^{\prime}(s)}\right),\pi_{v}=\arctan\!\left(\frac{-z_{t}^{\prime}(s)}{\sqrt{x_{t}^{\prime}(s)^{2}+y_{t}^{\prime}(s)^{2}}}\right).\end{aligned}$
@@
$\begin{aligned} \psi_{d}=\pi_{h}-\hat{\beta}_{c}-\arctan\!\left(\frac{y_{e}^{p}}{\Delta_{h}}\right),\theta_{d}=\pi_{v}+\hat{\alpha}_{c}+\arctan\!\left(\frac{z_{e}^{p}}{\Delta_{v}}\right),\end{aligned}$
@@
with positive adaptation gains $\gamma_{h},\gamma_{v}>0$, ensuring uniform semiglobal exponential stability of the cross-track errors for straight-line path following (Fossen and Aguiar, 2024).

### v2

@@ Abstract
This paper presents a fully data-driven control framework for autonomous underwater vehicles (AUVs) based on Data-Enabled Predictive Control (DeePC). The approach eliminates the need for explicit hydrodynamic modeling by exploiting measured input-output data to predict and optimize future system behavior. Classic DeePC was employed in the heading control, while a cascaded DeePC architecture is proposed for depth regulation. For 3-D waypoint path following, the Adaptive Line-of-Sight algorithm is extended to a predictive formulation and integrated with DeePC. All methods are validated in extensive simulation on the REMUS 100 AUV and compared with classical PI/PID control. The results demonstrate superior tracking performance and robustness of DeePC under ocean-current disturbances and nonlinear operating conditions, while significantly reducing modeling effort.
@@
Data-driven Model Predictive Control , Autonomous Underwater Vehicles , Cascaded DeePC , Predictive ALOS , 3-D path following
journal: Ocean Engineeringcodecodefootnotetext: Code available: https://github.com/SebsDevLab/DeePC_AUV.git
[…]
@@ 1 Introduction
Autonomous Underwater Vehicles (AUVs) have become central to modern marine exploration, inspection, and defense activities [16]. Ensuring reliable path-following and trajectory tracking in such missions requires controllers that can cope with nonlinear hydrodynamics and environmental disturbances [22]. These control objectives are typically achieved through a hierarchical control structure, which integrates a high-level guidance system responsible for generating appropriate reference signals, an intermediate-level controller tracking those references, and a low-level control system that directly manages the actuators [6]. A classic example of a high-level guidance system for AUVs is Line-of-Sight (LOS) guidance. The LOS method provides a geometric three-point guidance rule for steering the vehicle toward a target point on the desired path, effectively translating the spatial error into a heading command [6, 5]. More sophisticated LOS-based guidance laws have been summarized in [16], and the state-of-the-art lies in the Adaptive Line-of-Sight (ALOS) algorithm, which optimizes reference generation by compensating for varying environmental conditions [14]. In practice, these high-level guidance laws are commonly paired with reactive PI/PID controllers that do not account for future behavior or constraints. As a result, their performance deteriorates when faced with strong nonlinearities, actuator saturation, or delays in dynamics. Consequently, predictive control strategies have been introduced to AUVs, such as Model Predictive Control (MPC) [22, 21].
These approaches require a model of the AUV dynamics, which is complex and time-consuming [23]. The hydrodynamic coefficients that characterize added mass, drag, lift, and damping are often nonlinear, coupled, and highly dependent on factors such as vehicle geometry, speed, and surrounding flow conditions. Identifying these parameters requires extensive experiments that are time-consuming and context-specific [23], and the resulting models typically remain valid only within a limited operating range, reducing their applicability across different speeds, payloads, or environmental conditions. Such challenges have motivated data-driven control approaches, which exploit measured data to design and execute control laws without requiring an explicit parametric model of the system. Within this context, the Data-Enabled Predictive Control (DeePC) framework introduced in [7] has emerged as one of the most promising approaches. DeePC uses offline trajectories of inputs and outputs to solve a constrained optimization problem for optimal control inputs. This is particularly attractive for AUV applications, especially due to its potential to also handle complex and nonlinear dynamics [10].
@@
This paper is organized as follows. Notation is outlined in Section 2, and the general AUV dynamics, state-of-the-art, and theoretical foundations of DeePC are in Section 3, Section 4, and Section 5, respectively. Subsequently, the control architecture design and PALOS are described in Section 6-7, and numerical results obtained from high fidelity MSS-based simulations [12] are analyzed in Section 8. Additionally, the study includes stress tests addressing system nonlinearities and ocean currents to evaluate the robustness and identify the limitations of the proposed control approaches. The built-in PID controllers and the recently developed ALOS algorithm [14, 13] from the MSS toolbox are employed as benchmarks to contextualize the obtained results. Finally, the conclusion and future works are in Section 9.
@@ 2 Notation
To simplify the representation of the AUV dynamics, the state vector is expressed with respect to two coordinate systems: the North-East-Down (NED) frame and the body-fixed frame. Using the Euler angle transformation, the rotation matrix ${R}_{b}^{n}$ defined in (2) transforms velocities from the body-fixed frame $b$ to the NED frame $n$, as shown in (1[15], where $v_{n}$ and $v_{b}$ are velocities with respect to the NED and body-fixed frames, respectively. The inverse matrix ${R}_{n}^{b}=({R}_{b}^{n})^{-1}$ enables the transformation in the opposite direction. Here, $c\psi$ and $s\psi$ denote the cosine and sine of the angle $\psi$, respectively.
${v}_{n}={R}_{b}^{n}{v}_{b}$ (1)
$T_{k}=\begin{bmatrix}1&\sin\phi\tan\theta&\cos\phi\tan\theta\\[4.0pt] 0&\cos\phi&-\sin\phi\\[4.0pt] 0&\sin\phi/\cos\theta&\cos\phi/\cos\theta\end{bmatrix}$ (3)
@@ 3 Autonomous Underwater Vehicle
The AUV considered in this study is the REMUS 100, as implemented in Fossen’s MSS toolbox [12]. The model includes six degrees of freedom (6-DOF) dynamics with actuation provided by a three-bladed propeller, stern, and rudder fins, allowing control of surge, heave, and sway motions. The REMUS 100 is a representative small vehicle with a weight of approximately 31.9 kg, a length of 1.6 m, and a diameter of 0.19 m. [15] provides a detailed presentation of the mathematical models and theoretical concepts underlying longitudinal AUVs. In this paper, only the main relations relevant for control design and understanding the underlying nonlinearities are summarized in detail, while the specific REMUS 100 parameters are obtained from the MSS toolbox [12].
@@ 3.1 The rigid-body equations of an AUV
The dynamics of the AUV are represented using a standard 6-DOF rigid-body formulation [15]. The vehicle state $x=[\eta^{\top},\nu^{\top}]^{\top}$ is defined by its pose $\eta$ and velocity vector $\nu$. The pose consists of its position $x_{p},y_{p},z_{p}$ in the inertial NED frame and the attitude, consisting of the Euler angles $\phi,\theta,\psi$, which describe the orientation of the body frame relative to the NED frame in (6). The linear velocities $u$, $v$, $w$ (surge, sway, heave) and angular velocities $p$, $q$, $r$ (roll, pitch, yaw) in the velocity vector $\nu$ are expressed in the body frame given in (7). The relative velocity with respect to the surrounding water is given in (8), where $\nu_{c}=[u_{c},v_{c},w_{c},0,0,0]^{\top}$ contains the ocean currents expressed in the body frame.
@@
3.2 Parameter Introduction for AUV dynamic matrices
Prior to describing the components of the 6-DOF rigid-body equations of motion, some fundamental concepts regarding the vehicle geometry must be established. The REMUS 100 AUV is represented using the Myring hull equations [20] and, in the MSS toolbox, approximated as a spheroid [12]. The spheroid is defined by its longitudinal, transverse, and vertical semi-axes $a$, $b$, and $c$. Owing to the near-cylindrical shape of the vehicle, the transverse and vertical semi-axes are equal ($b=c$), and the geometry satisfies the relation in (11) in terms of the AUV’s length $L_{\mathrm{AUV}}$ and diameter $D_{\mathrm{AUV}}$. To ensure consistency with the vehicle mass, a rescaling of the semi-axes is performed according to (12), where $m$ is the vehicle mass and $\rho$ the fluid density. Solving these relations yields the semi-axis values $a$ and $b$ defining the equivalent spheroidal representation of the REMUS 100.
$\frac{L_{\mathrm{AUV}}^{2}}{a^{2}}+2\frac{D_{\mathrm{AUV}}^{2}}{b^{2}}=1$ (11)
$m=\frac{4}{3}\pi\rho ab^{2}$ (12)
$I_{x_{b}}=\frac{2}{5}m\,b^{2},$ (13)
$I_{y_{b}}=I_{z_{b}}=\frac{1}{5}m\,(a^{2}+b^{2}),$ (14)
$e=1-\left(\frac{b}{a}\right)^{2}$ (15)
$\beta_{0}=\frac{1}{e^{2}}-\frac{1-e^{2}}{2e^{3}}\ln\frac{1+e}{1-e}$ (17)
Following the $k$-factors introduced in [19], the added-mass contributions to the system inertia can be estimated. The factors are defined as
yielding the diagonal added-mass matrix
While classical hydrodynamics predicts zero added moment in roll (element $(4,4)$ of $M_{A}$) for a spheroidal hull [18], practical vehicles such as the REMUS 100 experience additional hydrodynamic contributions from the control surfaces and the propeller. Accordingly, this roll-related entry of the added-mass matrix is nonzero, $M_{A}(4,4)=r_{44}I_{x}$, capturing these effects in the simulation [12].
$U_{r}=\sqrt{u_{r}^{2}+v_{r}^{2}+w_{r}^{2}}$ (20)
$U_{rh}=\sqrt{u_{r}^{2}+v_{r}^{2}},\quad U_{rv}=\sqrt{u_{r}^{2}+w_{r}^{2}}$ (21)
3.3 The AUV dynamic matrices
In the MSS toolbox, $M_{A}$ is assumed constant, resulting in a constant $M$. The REMUS 100 exhibits both port–starboard and fore–aft symmetry, which simplifies the inertia representation according to [15] to the sparse structured form shown in (25).
@@ 3.3 The AUV dynamic matrices
[…]
$D(\nu_{r})=-\operatorname{diag}\Big\{\,X_{u}e^{-3U_{r}},\;Y_{v}e^{-3U_{r}},\;Z_{w},\;K_{p},\;M_{q},\;N_{r}\,\Big\}.$ (29)
@@
as well as lift forces in sway and heave, respectively,
$F_{\mathrm{lift}}=\tfrac{1}{2}\rho U_{r}^{2}\,S\,C_{L}(\alpha),\quad F_{\mathrm{drag}}=\tfrac{1}{2}\rho U_{r}^{2}\,S\,C_{D}(\alpha),$ (36)
$\tau_{LD,Z}$ $=-F_{\mathrm{drag}}\sin\alpha\;-\;F_{\mathrm{lift}}\cos\alpha.$
$\tau_{LD}=\left[\tau_{LD,X},0,\tau_{LD,Z},0,0,0\right]^{\top}$ (38)
The cross-flow drag forces and moments are modeled following the strip-theory formulation by [15] and [11]. The AUV hull is approximated as a series of two-dimensional strips of draft $T=D_{\mathrm{AUV}}$ and length $L=L_{\mathrm{AUV}}$, each experiencing a local flow velocity dependent on the sway and heave components $(v_{r},w_{r})$ and the corresponding angular rates $(r,q)$. The sectional drag is characterized by the two-dimensional drag coefficient $C_{\mathrm{d}}^{2\mathrm{D}}(v_{r},w_{r})$, yielding the cross-flow drag integrals in (39). The total generalized cross-flow drag vector acting on the vehicle is given in (40), representing the hydrodynamic forces and moments in sway, heave, pitch, and yaw due to lateral and vertical flow separation along the hull.
$\tau_{CF}=\left[0,\,Y_{CF},\,Z_{CF},\,0,\,M_{CF},\,N_{CF}\right]^{\top},$ (40)
@@ 3.4 AUV Forces and Moments
3.5 Ocean Current model
The current velocity expressed in the body-fixed frame is defined as
$\nu_{c}^{b}=\begin{bmatrix}u_{c},v_{c},w_{c}\end{bmatrix}^{\top}$. (45)
$\nu_{c}^{b}=\left(R_{b}^{n}\right)^{-1}{\nu}_{{c}}^{n}.$ (47)
$\dot{\nu}_{r}$ $=\dot{\nu}-\dot{\nu}_{c}$ (48)
$=M^{{-}1}\!\big(\tau\!+\!\tau_{LD}\!+\!\tau_{CF}{-}C(\nu_{r})\nu_{r}{-}D(\nu_{r})\nu_{r}{-}g(\eta)\big).$
@@ 4 Classic AUV Control
Since the control objective of this work is path-following, the propeller speed $n_{p}$ is fixed to a constant value corresponding to a nominal cruising velocity. The controller gains of the following PI/PID controllers were tuned by pole placement as described in [15] and [1].
@@ 4.2 Depth Control
$\theta_{d}$ $=K_{p_{z}}(z_{p}-z_{p,d})+K_{i_{z}}{\int_{0}^{t}(z_{p}-z_{p,d})\,\mathrm{d}\tau},$ (51)
$\delta_{s}$ $=-K_{p_{\theta}}(\theta-\theta_{d})-K_{d_{\theta}}q-K_{i_{\theta}}{\int_{0}^{t}(\theta-\theta_{d})\,\mathrm{d}\tau}.$
@@ 4.3 Adaptive Line-of-Sight Guidance Law
The ALOS guidance law [13] provides a 3-D formulation for path following of AUVs. It generates desired yaw and pitch angles, $\psi_{d}$ and $\theta_{d}$, to align the vehicle with a spatial path while adaptively compensating for unknown constant disturbances such as ocean currents. The spatial path is described by the position vector ${p}_{s}(s)=[x_{s}(s),\,y_{s}(s),\,z_{s}(s)]^{\top}$, parameterized by the path variable $s$ and contains all possible points on the path. This representation can encode smooth curves, splines, or straight lines. Its local orientation is given by the tangent vector ${t}_{t}(s)$, which in this work is simplified to a straight line between two consecutive waypoints as in (52).
@@
$\pi_{h}=\arctan\!\left(\frac{y_{t}^{\prime}(s)}{x_{t}^{\prime}(s)}\right),$ (53)
$\pi_{v}=\arctan\!\left(\frac{-z_{t}^{\prime}(s)}{\sqrt{x_{t}^{\prime}(s)^{2}+y_{t}^{\prime}(s)^{2}}}\right).$
@@
$\psi_{d}=\pi_{h}-\hat{\beta}_{c}-\arctan\!\left(\frac{y_{e}^{p}}{\Delta_{h}}\right),$ (55)
$\theta_{d}=\pi_{v}+\hat{\alpha}_{c}+\arctan\!\left(\frac{z_{e}^{p}}{\Delta_{v}}\right),$
@@
with positive adaptation gains $\gamma_{h},\gamma_{v}>0$, ensuring uniform semiglobal exponential stability of the cross-track errors for straight-line path following [13].